# AI-Powered CPS-Enabled Urban Transportation Digital Twin: Methods and Applications

### v2

@@ Abstract
We present methods and applications for the development of digital twins (DT) for urban traffic management. While the majority of studies on the DT focus on its “eyes,” which is the emerging sensing and perception like object detection and tracking, what really distinguishes the DT from a traditional simulator lies in its “brain,” the prediction and decision making capabilities of extracting patterns and making informed decisions from what has been seen and perceived. In order to add value to urban transportation management, DTs need to be powered by artificial intelligence and complement with low-latency high-bandwidth sensing and networking technologies, in other words, cyberphysical systems (CPS). We will first review the DT pipeline enabled by CPS and propose our DT architecture deployed on a real-world testbed in New York City. This paper can be a pointer to help researchers and practitioners identify challenges and opportunities for the development of DTs; a bridge to initiate conversations across disciplines; and a road map to exploiting potentials of DTs for diverse urban transportation applications.
Index Terms:
Digital twin, AI, Urban traffic management
I Introduction
@@
With recent explosive growth of literature on DTs, we would like to restrict the scope of this paper to applications in the urban setting, especially when vulnerable road users (VRU) (i.e., non-motorists such as pedestrian, bicyclists, other cyclists, or persons on personal conveyance [9]) are an integral part of the system and also potential users of the DT. We will primarily focus on use cases accounting for VRUs with improved traffic safety and efficiency.
The studies of DTs for urban traffic management, especially involving VRUs, are lacking, partly because the development of a DT for a system is non-trivial, particularly when involved with humans.
This paper presents methods and applications for the development of DTs for urban transportation systems. We will depict a DT pipeline prototype, leveraging the architecture of cyberphysical systems and AI methods. We will propose a DT for real-time traffic monitoring and optimization, based on an existing physical testbed deployed in New York City (NYC) leveraging cutting-edge sensing, communication, computing, and AI-based automation. The overall contributions of this paper include: (1) introducing AI methods used in the DT pipeline; (2) exploring the architecture of transportation DTs and propose a prototype for reference; and (3) identifying gaps and directions.
The rest of this paper is organized below. Section II introduces transportation DTs enabled by cyberphysical systems, and position this paper; Section III reviews the literature along the pipeline of a DT. Section IV demonstrates the architecture of our DT, building on a real-world testbed. Section V concludes our work, presents potential research directions and open questions.
II Preliminaries
@@
Cyberphyscial systems (CPS) [12] are smart systems that include engineered interacting networks of physical and computational components. CPS holds great potential to enable real-time applications thanks to emerging technologies in sensing, communication, and computing.
@@
A transportation CPS interlinks physical and cyber layers, where the cyber layer consists of sensing, networking, computing, and traffic management application modules (see Fig. 1). The DT encloses the cyber layer, and relies on all the modules for two-way interaction with the physical. To enable the technological development of a DT, a physical testbed is needed as a platform for sensing, computing, experimentation, evaluation, as well as design constraints determination. In Sec. III, we will review the CPS technological enablers needed for the development of a DT, and examine the testbed used for our proposed DT in Sec. IV.
We would like to stress that, this paper aims to discuss how AI algorithms and the architecture of CPS contribute to the urban T-DT (UT-DT) pipeline. In contrast, there are highly cited survey papers on T-DT, which are more focused on general transportation scenarios and applications, while AI might not be the key focus. Tab. I outlines the comparison of a partial set of related work.
TABLE I: Comparison of survey papers on DT
Ref. Topic Focus
[4][5][6] Comprehensive review for vehicle mobility apps Vehicular technology driven pipeline
[13] Comprehensive review for traffic safety and mobility CPS pipeline highlighting communication
[14] Comprehensive review for operation and maintenance apps Broad travel modes
Ours Semi-technical review and position for urban traffic apps AI-Powered CPS pipeline
@@
We first review various CPS and AI methods needed in the pipeline, and then present our T-DT instance. Accordingly, this paper is semi-survey, semi-technical. We employ such an organization, because we find that the publications on T-DT are normally segmented by different communities and journals, which could prevent researchers from understanding the entire pipeline, from upstream sensing and perception, to downstream transportation applications. For example, [15] on networking DT is primarily focused on the development of communication technology with evaluation on latency, even its use case is transportation. Transportation researchers, who hope to implement this system in real-world, have to seek more details about how traffic dynamics would impact the performance, which is unfortunately missing. This is likely because the authors belong to the society of communication and networking. Another example is that, [3] focused on vehicular DT, heavily rely on the foundational knowledge in CPS, which could be somewhat unfamiliar to many transportation researchers. Realizing such a gap in a large body of literature on T-DT, this paper aims to unify the knowledge by presenting a comprehensive summary upfront, and exemplify the pipeline following the summary. After all, the development of a T-DT calls for the interdisciplinary collaboration across transportation, electrical engineering, mechanical engineering, computer science, and human-machine interaction. To foster the readability of the paper, in the next section, we first offer a comprehensive state-of-the-art review on related work along the CPS pipeline.
III Related work on CPS-enabled DT
@@
III-A Sensing and perception
Sensors are the “eyes” (and “ears”) of a DT. Tab. II summarizes the pros and cons of each sensing technology, namely, mobile devices, on-board vehicles, and roadside infrastructure, for urban traffic applications. The former two are mobile sensors with wider spatial coverage but challenge in precision because of moving references, while the latter at fixed locations could face limited sensing ranges and coverage.
TABLE II: Sensors for urban traffic applications (partially adapted from [16, 17])
@@ III-A Sensing and perception
III-B Object detection and tracking
Object detection identifies and classifies objects within the environment using sensors like cameras, Lidar, and Radar. Object tracking is the process of monitoring the detected objects over time to determine their position and movement. Multi-object tracking is concerned with maintaining the identity of the objects and generating their trajectories. Trajectory prediction involves forecasting the future paths of detected and tracked objects. These tasks highly rely on training datasets for urban traffic scenes. Note that there is a much larger size of public datasets collected from on-board vehicles [25], but fewer from other sensor types. Here we thus summarize commonly used and emerging datasets from non-vehicle sensors in Tab. III.
TABLE III: Public datasets for urban traffic scenes (On-board vehicle datasets can be found in [25].)
@@ III-B Object detection and tracking
Predicting future trajectories of detected objects is often a crucial part of safety-critical applications. Numerous deep neural network models have emerged as competitive candidates for trajectory prediction over the past few years. Majority of modern architectures for predicting future trajectories of detected objects adopt Recurrent Neural Networks which is responsible for predicting future object positions based on their historical coordinates, together with generative components which handles the variation and flexibility in social interactions [78, 79]. Neural Social Physics model [80] incorporates learnable parameters into explicit physics models built on top of neural networks. SemanticFormer [81] seeks more structural and humanized environment understanding by constructing a semantic knowledge graph. TrajNet++ [82], TDOR [83], and CASPNet++ [84] predict the distributions of future trajectories based on occupancy grid maps. Models like FRM [85] and PPT [86] decompose the prediction task by taking a multi-stage approach. Larger amounts of data comprised of multiple modalities and more comprehensive frameworks have shown increasing importance as is demonstrated by UniTraj++ [87]. Unlike the tracking algorithms, specific training or fine-tuning is often required before the deployment of trajectory forecasting models to unseen scenarios.
III-C Real-time video analytics
Developing end-to-end real-time video analytics systems on a large scale for time-sensitive and safety-critical traffic and crowd management applications presents challenges. Video analytics requires the collection and processing of large volumes of video data, which can be resource-intensive and costly. Optimizing computation and network resource usage while maintaining or enhancing the accuracy of analytical results can be challenging. This challenge is further complicated by the need to adapt to varying network conditions, computational resources, and dynamic scene changes in real-time. Tab. IV provides a comparative analysis of various approaches to address these challenges. The approaches differ in their focus–some prioritize reducing latency and resource consumption, while others emphasize maintaining or enhancing accuracy, especially under constrained conditions. This comparison provided in Tab. IV highlights the trade-offs inherent in real-time video analytics and emphasizes different optimization methods to balance throughput, accuracy, energy consumption, and computational efficiency across diverse deployment scenarios, including edge devices, cloud platforms, and hybrid environments.
TABLE IV: Comparison of various approaches and optimization objectives in video analytics.
@@ III-C Real-time video analytics
III-D Communication and networking
A DT for safety-critical applications requires real time communications with aggressively low latency. We explore issues related to low latency communications, and survey component technologies and protocols that can be utilized to achieve very low latency.
III-D1 Real-time requirements and low latency targets
Sensor and control data in a real time system is subject to latency created by the stages, namely, (1) data acquisition from traffic participants (such as camera recordings and encoding, harvesting data from autonomous vehicles, and collect information from fiber); (2) transmission of data across communications links from sensors to inferencing servers using communications protocols such as Transmission Control Protocol (TCP), Unreliable Data Protocol (UDP) and Real-Time Streaming Protocol (RTSP); (3) data preprocessing (video decoding, and cropping); (4) AI inferencing; (5) higher level reasoning about required feedback to traffic participants; and (6) sending feedback to traffic participants across communications links via low-latency broadcast or dedicated channels over wired and wireless.
@@ III-D1 Real-time requirements and low latency targets
III-D2 Communication techniques and protocols
@@
Tab. V summarizes the key characteristics of commonly used IoT communication protocols, Message Queuing Telemetry Transport (MQTT) [120], Constrained Application Protocol (CoAP) [121], and Hypertext Transfer Protocol (HTTP) [122]. Among them, MQTT emerges as a practical choice, due to its combination of low latency, high scalability, and reliable delivery mechanisms. Its lightweight publish/subscribe model, Quality of Service (QoS) guarantees, and session management make it well suited for real-time data exchange in dynamic, safety-critical environments like urban transportation systems [123, 124].
TABLE V: Communication protocol comparison for real-time digital twin systems.
@@
IV Proposed DT pipeline
Figure 2: Architecture of the proposed DT pipeline: Urban Transportation DT (UT-DT).
In this section, we present a DT architecture for UT-DT, based on the sensing/communication/computing testbed deployed in NYC (Fig. 2). The proposed architecture is enabled by cameras and LiDARs, high speed communications, and edge cloud computing.
Figure 3: Diagram of intersection safety warning use case.
IV-A Physical Infrastructure
Pilot experiments are executed at the signalized intersection of Amsterdam Avenue (major) and 120th Street (minor) near Columbia campus in NYC. The road geometry and traffic statistics are summarized in Fig. 4.
Figure 4: Traffic statistics for NYC’s intersection of 120th St. & Amsterdam Ave.
IV-A1 Technological enabler
The physical functionality of the proposed UT-DT is built on top of the COSMOS testbed (“Cloud enhanced Open Software defined mobile wireless testbed for city-Scale deployment”) [125], developed for real-world research, development, and deployment of city-scale advanced wireless communications and innovative applications [125]. It targets the technology “sweet spot” of ultra-high bandwidth and ultra-low latency, a capability that will enable a broad class of safety-time critical applications. Deployed in West Harlem, NYC, next to Columbia University campus, the COSMOS testbed is an enabler for research on the design, development, and deployment of a DT for urban traffic management.
At the intersection of Amsterdam Avenue and 120th Street, cameras, LiDAR, and wireless sensing and communication nodes are deployed. The cameras employ H264 encoding with an I-frame interval set to 10 frames. The live RTSP video streams–with 4K resolution at 30 fps–were processed on a COSMOS edge server equipped with an A100 GPU. The LiDAR is a Velodyne VLP-32C with up to 200m range.
IV-B Digital twin
Building on data collected from the physical world, we built an Unreal Engine based simulation platform for generating photorealistic street scenes. Moving objects are populated into an integrated SUMO-CARLA simulation platform to validate our safety warning application [126], where vehicle locations are synchronized from the real-world MQTT messages.
[…]
(a) COSMOS in CARLA
(b) COSMOS in SUMO
(c) Boundless with snow
(d) Boundless with rain
Figure 5: (Top) Carla-SUMO simulator; (Bottom) Qualitative examples from Boundless. Bounding boxes for vehicles are shown in blue, and pedestrians are shown in green.
IV-C Use cases
@@
The value of a DT for urban transportation management ranges from traffic state prediction, spatiotemporal traffic flow forecasting, to urban planning and policy making. Here we focus on two that are particularly important in urban settings, leveraging video analytics.
IV-C1 Use case 1: Micromobility detection
In the U.S., the e-bike had a market size of almost \$2 billion in 2022, with a projected growth of 15.6% from 2023 to 2030. In 2022, 1.1 million e-bikes were sold, four times as many as were sold in 2019. In 2019, 136 million trips were made using micromobility, a 60% increase from 2018. Because their presence and surge have profound implications for road safety and urban traffic management, automatic detection of micromobility for regulation becomes increasingly important. However, there does not exist a standard AI model nor a dataset for the object detection of this emerging transportation mode.
AI model evaluation
[…]
TABLE VI: Class-wise recall (%) on the Micromobility test set. Higher values indicate better detection performance.
Model Bicycle $\uparrow$ e-Bike $\uparrow$ Motorcycle $\uparrow$
YOLOv9e 52.9 54.7
Ours 68.1 61.0 91.8
Figure 6: Four sample frames showing the small scale of micromobility instances from infrastructure-based 4K resolution cameras in the COSMOS testbed.
IV-C2 Use case 2: VRU safety warning
@@
Safety-critical applications, making automated life-and-death decisions such as collision avoidance warning between automobiles and pedestrians, need to activate at the precise time and the right moment with bounded latency. Safety-critical systems are often time-critical and reactive, because they need to react to external signals in precise time [127], and require tasks to be executed in a timely manner. Theses systems hold the substantial potential to enhance road safety and save lives. They are, however, risky to test and run, because rare events like automobile collisions are challenging and unethical to replicate. Thanks to emerging technologies in ubiquitous sensing, low-latency high-bandwidth communication, high-speed computing and AI, safety critical applications could potentially be modeled, simulated, processed, and tested in a DT. The applications of DTs on safety-critical scenarios, however, remains understudied, because it necessities short runtime and real-time reaction, posing high requirements for communication and computing technologies, pipeline architecture design, and testing. We summarize the safety-critical applications to address conflict risks between vehicles and VRUs in Tab. VII, and offer an outlook of safety guarantee related methods in Sec. V-A3.
TABLE VII: Literature review on safety warning to VRUs.

### v3

# AI-Powered CPS-Enabled Vulnerable-User-Aware Urban Transportation Digital Twin: Methods and Applications

@@ Abstract
We present methods and applications for the development of digital twins (DT) for urban traffic management. While the majority of studies on the DT focus on its “eyes,” which is the emerging sensing and perception like object detection and tracking, what really distinguishes the DT from a traditional simulator lies in its “brain,” the prediction and decision making capabilities of extracting patterns and making informed decisions from what has been seen and perceived. In order to add value to urban transportation management, DTs need to be powered by artificial intelligence and complement with low-latency high-bandwidth sensing and networking technologies, in other words, cyberphysical systems. This paper can be a pointer to help researchers and practitioners identify challenges and opportunities for the development of DTs; a bridge to initiate conversations across disciplines; and a road map to exploiting potentials of DTs for diverse urban transportation applications.
footnotetext: This work has been published in IEEE Transactions on Intelligent Transportation Systems (IEEE T-ITS).footnotetext: Yongjie Fu, Mehmet K.Turkcan, Zhaobin Mo, and Xuan Di are with the Department of Civil Engineering and Engineering Mechanics, Columbia University. Emails: {yf2578, mkt2126, zm2302, sharon.di}@columbia.edufootnotetext: Mehmet K. Turkcan and Mahshid Ghasemi contributed equally to this work.footnotetext: Xuan Di, Zoran Kostic, and Gil Zussman are also with the Data Science Institute, Columbia University.footnotetext: Mahshid Ghasemi, Chengbo Zang, Abhishek Adhikari, Zoran Kostic, and Gil Zussman are with the Department of Electrical Engineering, Columbia University. Emails: {mg4089, cz2678, abhishek.adhikari, zk2172, gil.zussman}@columbia.edu
Keywords: Digital twin, AI, Urban traffic management
1 Introduction
@@
With recent explosive growth of literature on DTs, we would like to restrict the scope of this paper to applications in the urban setting, especially when vulnerable road users (VRU) (i.e., non-motorists such as pedestrian, bicyclists, other cyclists, or persons on personal conveyance [199]) are an integral part of the system and also potential users of the DT. The studies of DTs for urban traffic management, especially involving VRUs, are lacking, partly because the development of a DT for a system is non-trivial, particularly when involved with humans. This paper presents methods and applications for the development of DTs for urban transportation systems. We will depict a DT pipeline prototype, leveraging the architecture of cyberphysical systems and AI methods.
The rest of this paper is organized below. Section 2 introduces transportation DTs enabled by cyberphysical systems, and position this paper; Sections 3 and 4 review each component within CPS, and use cases for urban transportation, respectively. Section 5 demonstrates the available testbeds for DT testing and validation. Section 6 concludes our work, presents potential research directions and open questions.
2 Preliminaries
@@
Cyber-Physical Systems (CPS) [75] are smart systems that include engineered interacting networks of physical and computational components. CPS holds great potential to enable real-time applications thanks to emerging technologies in sensing, communication, and computing.
@@
A transportation CPS interlinks physical and cyber layers, where the cyber layer consists of sensing, networking, computing, and traffic management application modules (see Fig. 1). The DT encloses the cyber layer, and relies on all the modules for two-way interaction with the physical. To enable the technological development of a DT, a physical testbed is needed as a platform for sensing, computing, experimentation, evaluation, as well as design constraints determination. Tab. 1 outlines the comparison of a partial set of related work.
Table 1: Comparison of relevant review papers on DT
Ref Topic Focus Urban Applications CPS role in DT VRU Scope Real Testbeds Integration
Ours Review and position for urban transportation AI-powered CPS architecture Intersection safety warning; Signal control Interlinks physical and cyber layers, DT architecture VRU safety and efficiency A summary of relevant physical testbeds Bidirectional: data exchange, feedback control
[179] Overview for vehicle mobility Connected and automated vehicles (CAV) CAV design and testing (ADAS, ADS) Cloud-based CPS as methodologies Pedestrian safety in CAV testing Mcity, GoMentum Hardware, software, model-in-the-loop
[215] Review for vehicle mobility, architecture, case studies AI-based data-driven cloud-edge-device pipeline Personalized ACC, traffic management, driver modeling DT conceptualized as higher-level evolution of CPS Modeled in physical layer of DT Highway 237 in Mountain View, CA Data and service integration across twins
[89] Review for traffic safety and mobility Hierarchical architecture, cybersecurity Broad transportation safety and mobility services Enables feedback between physical and cyber components As mobile nodes in physical layer, modeled as individual asset for behavior analysis Testbeds mentioned for some literature, but not systematic Integration of data from different levels of twins, and with commun. gateway
[87] Review for smart cities applications Enabling technologies: ML, IoT, CPS, blockchain Transport, water management and automotive technology Facilitates data collection and decision making, DT architecture Ped. safety in Ford’s DT for predictive headlights Durham County Council’s streetlights, IIT’s sewage treatment plant, Insurance Institute for Highway Safety test track Interoperability among different DTs within the smart city ecosystem
[219] Review for transport operation and maintenance Varied levels of adoption across different transport sectors Broad intermodal and ground travel modes CPS enables DT through sensing communication control loop Pedestrian traffic management at train stations N/A Real-time data exchange and synchronization of decision making, integration across platforms
[223] Review for the full lifecycle of transport infrastructure Functions, architectures, integration along the lifecycle Planning and construction, operation and maintenance, and decommissioning and renewal N/A Integrated into an interoperable system for pedestrian pathways, infrastructure, micro-traffic behaviors N/A Data and system integration across various lifecycle stages
@@
The publications on T-DT are normally segmented by different communities and journals, which could prevent researchers from understanding the entire pipeline, from upstream sensing and perception, to downstream transportation applications. For example, [136] on networking DT is primarily focused on the development of communication technology with evaluation on latency, even its use case is transportation. Transportation researchers, who hope to implement this system in real-world, have to seek more details about how traffic dynamics would impact the performance, which is unfortunately missing. This is likely because the authors belong to the society of communication and networking. Another example is that, [118] focused on vehicular DT, heavily rely on the foundational knowledge in CPS, which could be somewhat unfamiliar to many transportation researchers. Realizing such a gap in a large body of literature on T-DT, this paper aims to unify the knowledge by presenting a comprehensive summary upfront, and exemplify the pipeline following the summary with selected applications. After all, the development of a T-DT calls for the interdisciplinary collaboration across transportation, electrical engineering, mechanical engineering, computer science, and human-machine interaction. To foster the readability of the paper, in the next section, we first offer a comprehensive state-of-the-art review on related work along the CPS pipeline.
3 CPS-enabled DT
@@
3.1 Sensing and perception
Sensors are the “eyes” (and “ears”) of a DT. Tab. 2 summarizes the pros and cons of each sensing technology, namely, mobile devices, on-board vehicles, and roadside infrastructure, for urban traffic applications. The former two are mobile sensors with wider spatial coverage but challenge in precision because of moving references, while the latter at fixed locations could face limited sensing ranges and coverage.
Table 2: Sensors for urban traffic applications (partially adapted from [200, 256])
@@ 3.1 Sensing and perception
3.2 Object detection and tracking
Object detection identifies and classifies objects within the environment using sensors like cameras, Lidar, and Radar. Object tracking is the process of monitoring the detected objects over time to determine their position and movement. Multi-object tracking is concerned with maintaining the identity of the objects and generating their trajectories. Trajectory prediction involves forecasting the future paths of detected and tracked objects. These tasks highly rely on training datasets for urban traffic scenes. Note that there is a much larger size of public datasets collected from on-board vehicles [45], but fewer from other sensor types. Here we thus summarize commonly used and emerging datasets from non-vehicle sensors in Tab. 3.
Table 3: Public datasets for urban traffic scenes (On-board vehicle datasets can be found in [45].)
@@ 3.2 Object detection and tracking
Predicting future trajectories of detected objects is often a crucial part of safety-critical applications. Numerous deep neural network models have emerged as competitive candidates for trajectory prediction over the past few years. Majority of modern architectures for predicting future trajectories of detected objects adopt Recurrent Neural Networks which is responsible for predicting future object positions based on their historical coordinates, together with generative components which handles the variation and flexibility in social interactions [5, 81]. Neural Social Physics model [246] incorporates learnable parameters into explicit physics models built on top of neural networks. SemanticFormer [190] seeks more structural and humanized environment understanding by constructing a semantic knowledge graph. TrajNet++ [100], TDOR [80], and CASPNet++ [schäfer2023caspnet] predict the distributions of future trajectories based on occupancy grid maps. Models like FRM [160] and PPT [122] decompose the prediction task by taking a multi-stage approach. Larger amounts of data comprised of multiple modalities and more comprehensive frameworks have shown increasing importance as is demonstrated by UniTraj++ [57]. Unlike the tracking algorithms, specific training or fine-tuning is often required before the deployment of trajectory forecasting models to unseen scenarios.
3.3 Real-time video analytics
Developing end-to-end real-time video analytics systems on a large scale for time-sensitive and safety-critical traffic and crowd management applications presents challenges. Video analytics requires the collection and processing of large volumes of video data, which can be resource-intensive and costly. Optimizing computation and network resource usage while maintaining or enhancing the accuracy of analytical results can be challenging. This challenge is further complicated by the need to adapt to varying network conditions, computational resources, and dynamic scene changes in real-time. Tab. 4 provides a comparative analysis of various approaches to address these challenges. The approaches differ in their focus–some prioritize reducing latency and resource consumption, while others emphasize maintaining or enhancing accuracy, especially under constrained conditions. This comparison provided in Tab. 4 highlights the trade-offs inherent in real-time video analytics and emphasizes different optimization methods to balance throughput, accuracy, energy consumption, and computational efficiency across diverse deployment scenarios, including edge devices, cloud platforms, and hybrid environments.
Table 4: Comparison of various approaches and optimization objectives in video analytics.
@@ 3.3 Real-time video analytics
3.4 Communication and networking
A DT for safety-critical applications requires real-time communications with aggressively low latency. We explore issues related to low latency communications, and survey component technologies and protocols that can be utilized to achieve very low latency.
3.4.1 Real-time requirements and low latency targets
Sensor and control data in a real-time system is subject to latency created by the stages, namely, (1) data acquisition from traffic participants (such as camera recordings and encoding, harvesting data from autonomous vehicles, and collect information from fiber); (2) transmission of data across communications links from sensors to inferencing servers using communications protocols such as Transmission Control Protocol (TCP), Unreliable Data Protocol (UDP) and Real-Time Streaming Protocol (RTSP); (3) data preprocessing (video decoding, and cropping); (4) AI inferencing; (5) higher level reasoning about required feedback to traffic participants; and (6) sending feedback to traffic participants across communications links via low-latency broadcast or dedicated channels over wired and wireless.
@@ 3.4.1 Real-time requirements and low latency targets
3.4.2 Communication techniques and protocols
@@
Tab. 5 summarizes the key characteristics of commonly used IoT communication protocols, Message Queuing Telemetry Transport (MQTT) [161], Constrained Application Protocol (CoAP) [181], and Hypertext Transfer Protocol (HTTP) [60]. Among them, MQTT emerges as a practical choice, due to its combination of low latency, high scalability, and reliable delivery mechanisms. Its lightweight publish/subscribe model, Quality of Service (QoS) guarantees, and session management make it well suited for real-time data exchange in dynamic, safety-critical environments like urban transportation systems [12, 20].
Table 5: Communication protocol comparison for real-time DT systems.
@@
4 Use cases for urban transportation
@@
The value of a DT for urban transportation management ranges from traffic state prediction, spatiotemporal traffic flow forecasting, to urban planning and policy making. Here we primarily focus on those that are particularly important in urban settings leveraging video analytics, namely, VRU safety warning at road intersections and traffic signal control.
4.1 Safety-time critical applications
@@
Safety-critical applications [6], making automated life-and-death decisions such as collision avoidance warning between automobiles and pedestrians, need to activate at the precise time and the right moment with bounded latency. Exemplary systems include autonomous driving, Industry 5.0, and mobile and collaborative robots, strictly entailing precise time, computing, communication with bounded latency and AI [23]. Theses systems hold the substantial potential to enhance road safety and save lives. They are, however, risky to test and run, because rare events like automobile collisions are challenging and unethical to replicate. Thanks to emerging technologies in ubiquitous sensing, low-latency high-bandwidth communication, high-speed computing and AI, safety critical applications could potentially be modeled, simulated, processed, and tested in a DT. The applications of DTs on safety-critical scenarios, however, remains understudied, because it necessities short runtime and real-time reaction, posing high requirements for communication and computing technologies, pipeline architecture design, and testing.
Intersections, where sixty percent of crashes happen [88, 59], are critical bottlenecks of an urban transportation network. To improve urban road safety and increase traffic capacities, safety warning is the key. We summarize the safety-critical applications to address conflict risks between vehicles and VRUs in Tab. 6, and offer an outlook of safety guarantee related methods in Sec. 6.1.3.
Table 6: Literature review on safety warning to VRUs.